OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 5(x + 2)>= 7x + 2(x - 1)? Answer 3 >= x x <= 3 A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left.

OpenStudy (anonymous):

\[5x+10>=7x+2x-2\] \[5x-9x>=-2-10\] \[-4x>=-12\] \[x<=3\]